A cart is moving at 2 m/s then accelerates to 10 m/s in 4 seconds. Calculate the cart's acceleration.


10 m/s2


2.5 m/s2


4 m/s2


2 m/s2


1
Expert's answer
2021-03-24T19:39:56-0400
"a=\dfrac{v-v_0}{t}=\dfrac{10\ \dfrac{m}{s}-2\ \dfrac{m}{s}}{4\ s}=2\ \dfrac{m}{s^2}."